A copper wire has a diameter of 4.00mm. If the drift speed of the electrons in the wire is 0.0002 m/s, calculate the current flowing in the wire. Assume each copper atom contributes two free electrons to the current. The density of copper is 8.93 g/cm^3 and it's molar mass is 63.5 g/mol.
1
Expert's answer
2020-10-05T13:32:58-0400

number density of electron(n)-

n = "\frac{\rho N_a}{M}"

Here, "\rho" is the density of material in g/m3, Na is Avagadro Number, M is the molecular mass in g/mol.

n = "\frac{8930000 \times 6.022 \times 10^{23}}{63.5}"

= 8.5 "\times" 1028 atoms per m3

Current in the conductor (I)-

I = neAvD

Here, e is the electronic charge in C, A is the cross-section area of wire in m2, vD is the drift velocity of electron in m/s

I = (2 "\times" 8.5 "\times" 1028)(1.6 "\times" 10-19)("\pi \times" 4 "\times 10^{-6}" )(0.0002)

= 68.36 A